### v1

# Low-hallucination Synthetic Captions for Large-Scale Vision-Language Model Pre-training

Abstract
In recent years, the field of vision-language model pre-training has experienced rapid advancements, driven primarily by the continuous enhancement of textual capabilities in large language models. However, existing training paradigms for multimodal large language models heavily rely on high-quality image-text pairs. As models and data scales grow exponentially, the availability of such meticulously curated data has become increasingly scarce and saturated, thereby severely limiting further advancements in this domain. This study investigates scalable caption generation techniques for vision-language model pre-training and demonstrates that large-scale low-hallucination synthetic captions can serve dual purposes: 1) acting as a viable alternative to real-world data for pre-training paradigms and 2) achieving superior performance enhancement when integrated into vision-language models through empirical validation. This paper presents three key contributions: 1) a novel pipeline for generating high-quality, low-hallucination, and knowledge-rich synthetic captions. Our continuous DPO methodology yields remarkable results in reducing hallucinations. Specifically, the non-hallucination caption rate on a held-out test set increases from $48.2\%$ to $77.9\%$ for a 7B-size model. 2) Comprehensive empirical validation reveals that our synthetic captions confer superior pre-training advantages over their counterparts. Across 15 vision language tasks, the model trained with our data achieves a significant performance gain of at least $6.2\%$ compared to alt-text pairs and other previous work. In 20 common cognitive domains, the model trained with our data outperforms other counterparts by at least $6.8\%$. Meanwhile, it also offers considerable support in the text-to-image domain. With the support of our dataset, the FID score is reduced by $17.1$ on a real-world validation benchmark and $13.3$ on the MSCOCO validation benchmark. 3) We will release Hunyuan-Recap100M, a low-hallucination and knowledge-intensive synthetic caption dataset, which will enable breakthroughs in vision language pre-training, cross-modal generation, and related applications.
1 Introduction
With the enormous power of large language models, remarkable performance gains have recently been achieved in a variety of tasks in natural language processing (NLP), computer vision (CV), and also in cross-modal fields (Brown et al., 2020; Chung et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a; Dosovitskiy et al., 2021; Alayrac et al., 2022; Liu et al., 2023; Achiam et al., 2023; Wang et al., 2024b; Hurst et al., 2024). The Large Language Model (LLM) and the Vision Language Model (VLM) are usually equipped with Transformer (Vaswani et al., 2017) as the backbone and then pre-trained with a tremendous amount of unlabeled data. The strong representation ability of the model, the massive amount of data, and the effective means of training make the foundation models powerful for successfully solving the tasks of vision and language.
While text data has experienced exponential growth and fuels large language models, vision language model pre-training confronts a unique data bottleneck: high-quality text-image alignments remain scarce despite their critical role in training vision language models. Most traditional approaches to building large-scale multimodal datasets mainly rely on Web-based alt-text extraction (Sharma et al., 2018; Changpinyo et al., 2021; Ordonez et al., 2011; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Gadre et al., 2023; Liu et al., 2022; Gu et al., 2022). The web-based alt-text data typically undergoes a lengthy processing pipeline, involving stages such as text quality cleaning, image low-quality filtering, and cross-modal relevance validation, ultimately producing image-text pairs. Although significant data loss occurs at each processing stage and even with these efforts, the final output still suffers from inherent limitations including a low text informativeness density and partial relevance deficiencies between visual content and the corresponding descriptions. As the naturally existing image-text pair data is exhausted, and the quality of data produced by traditional image-text extraction methods drops significantly as the scale expands, the marginal effect they generate in the training of VLMs greatly diminishes. Therefore, new methods are needed to synthesize large-scale, high-quality multimodal data.
Recently, several exploratory investigations have been carried out in multi-modal synthetic data, which have, to some extent, demonstrated their positive impact on various cross-modal tasks (Li et al., 2023a; Betker et al., 2023; Yu et al., 2024; Li et al., 2024b). However, there are significant defects in all of them, mainly reflected in the presence of many hallucinations or a relatively low level of information content. As shown in Figure 1, there are a large number of hallucinations (labeled in red) in the previous synthetic captions, and the captions are rather monotonous, lacking informative expressions. Therefore, in this paper, we focus on addressing two issues: 1. How to establish an effective pipeline for generating low-hallucination synthetic captions? 2. How can more valuable information be injected into synthetic captions to increase the information density of synthetic data?
Figure 1: Examples of the original caption, previous synthetic caption and our synthetic caption.
How to establish an effective pipeline for generating low-hallucination synthetic captions? Large-scale synthetic captions inherently depend on robust tools, and vision language models emerge as a natural and fitting choice. However, hallucination is an inherent problem in the multimodal large language model (Bai et al., 2024). As shown in Figure 1, the data in the Recap-DataComp-1B (Li et al., 2024b) dataset contains a significant amount of hallucinatory information. We further demonstrate the harmfulness of a large amount of hallucinatory data in Section 4. The inevitable and most crucial issue in caption synthesis lies in resolving the hallucination problem. Existing work (Bai et al., 2024) has discussed the causes of hallucinations from the perspectives of data, model, training, and inference. During generation, as the length of the sequence grows, self-attention will focus more on previously generated text tokens, i.e., the attention to visual content is diluted(Wang et al., 2023b, a, 2024a; Huang et al., 2024). We agree that hallucinations in VLMs are mainly caused by uncertainty in decoding visual-related intentions. Therefore, we propose the use of Direct Preference Optimization (DPO) method to eliminate hallucinations in synthesized captions by increasing the confidences of real visual contents and suppressing the confidences of hallucinatory visual contents. Fortunately, this method has proven to be highly effective. Furthermore, in the synthetic caption task, we reveal that DPO adheres to a significant scaling law within a specific data size range. Upon saturation, resampling can further elevate the quality of synthetic captions. As presented in Table 1, when applied to the same dataset, our method significantly enhances the non-hallucination rate, elevating it from 29.7% to 73.3%, in contrast to the approach used in Recap-DataComp-1B (Li et al., 2024b). We elaborate on our hallucination evaluation metrics in Section 4.1 and the continuous DPO method in Section 3.
Dataset # Length # Details $\uparrow$ # None Hallucinations $\uparrow$ # Hallucination Rate $\downarrow$
Capfusion-120M 23.5 2.58 72.3% 14%
Our Method 80.4 6.9 73.9% 5.4%
Recap-DataComp-1B 49.41 6.95 29.7% 24.9%
Our Method 72.60 6.64 73.3% 5.64%
Hunyuan-Recap100M 103.15 8.1 77.9% 4.2%
Table 1: A comparison between existing synthetic caption datasets and our proposed dataset in terms of the number of hallucinations and the hallucination rate. Here, "Length" represents the average length of the caption, "Details" represents the average number of visual details that appear in each caption, "None Hallucinations" indicates the proportion of non-hallucinatory data in the overall data, and "hallucination rate" represents the proportion of hallucinations among all details. A total of 1,000 data samples were taken, and our method reports the corresponding results on the respective data with our synthetic caption pipeline respectively.
How can more valuable information be injected into synthetic captions to increase the information density of synthetic data? The answer to this question can be rather straightforward: infuse as much knowledge as possible into the caption via a well-trained VLM. The injection of knowledge into synthetic captions can be derived from two primary sources. The first source is external text, such as alt-text, that is relevant to the image, while the second source stems from world knowledge encapsulated within a VLM. Usually, a VLM model tends to provide captions that are vague but correct. Therefore, we need to train a VLM model so that it can fully integrate the knowledge in text and its own world knowledge, ultimately generating synthetic captions with high informational value. Specifically, in our methodology, we employ a powerful VLM, namely GPT-4o, in conjunction with a meticulously crafted prompt that encompasses detailed requirements. This combination is utilized to generate labeled data, with a particular emphasis on extracting and focusing on the key information embedded within alt-text. For instance, we focus more on the name of a specific entity, using oxalic acid as an illustration here. After that, we perform manually post-processing on obvious erroneous data to ensure that the labeled data is accurate, contains the key information, and excludes the noise in the injected information. Finally, we train an open source VLM model, specifically Qwen2-VL-7B, using the training data. For details of the prompt and data processing, see Section 3.
By comprehensively leveraging the aforementioned methods, we train a recaption model and utilize it to generate 100 million pieces of data. Our model has achieved a notable enhancement in the evaluation metrics for low-hallucination quantification. Moreover, we carried out a verification of the data’s validity. The results show that vision-language models pre-trained with Hunyuan-Recap100M exhibit substantially improved overall performance compared to those using alt-text data or conventional synthetic captions, particularly showing enhanced capabilities in perception tasks.
Our contributions are as follows.
(1) We tackle the challenge of constructing a synthetic caption pipeline capable of generating captions that are both low in hallucination and rich in knowledge, thereby alleviating the scarcity of high-quality multimodal data essential for vision-language model pre-training.
(2) We propose a series of effective methods, knowledge-enriching SFT and continuous DPO, to eliminate hallucinations and inject knowledge into synthetic captions. Eventually, we trained a recaption model, whose proportion of hallucination-free results far exceeds that of the existing best models. The proportion of hallucination-free results has increased from $48.2\%$ to $77.9\%$. It also exhibits favorable performance in knowledge diversity, as reflected in a more accurate cognition of real-world entities.
(3) We will release a high-quality synthetic caption dataset comprising $100M$ image-text pairs, Hunyuan-Recap100M, with the aim of fostering further advancements within the research community in areas like vision language model pre-training, text-to-image, and text-to-video generation.
2 Related Work
Vision-Language Model. Vision Language Models (VLMs) can be classified into two distinct categories. The first category consists of models that rely on dual encoders, specifically a vision encoder and a text encoder. These models are typically trained in a contrastive fashion and exhibit remarkable proficiency in tasks such as image recognition. Prominent examples within this group include CLIP (Radford et al., 2021), FLIP (Li et al., 2023b) and SigLIP (Zhai et al., 2023b). The second category of methods focuses on aligning the visual modality with a Large Language Model (LLM). With the rapid development of LLM, the astonishing success of GPT (Brown et al., 2020; OpenAI, 2023), LLaMA (Touvron et al., 2023a, b; Grattafiori et al., 2024), and DeepSeek-LLM (Liu et al., 2024; Guo et al., 2025) has made the empowerment of large language models with visual perception an important and popular research direction. Flamingo (Alayrac et al., 2022) is the first large-scale vision language model, which aligns the visual encoder with a large language model. LLaVA (Liu et al., 2023) presents an instruction tuning methodology for training large multimodal models, thus charting the course for VLM training. Subsequent efforts have been made to refine the details of each module and continuously optimize the overall performance of the model (Wang et al., 2024b; Wu et al., 2024; McKinzie et al., 2024; Chen et al., 2024b; Liu et al., 2024). Although, compared with former multimodal models (Zeng et al., 2021; Singh et al., 2021; Diao et al., 2022; Wang et al., 2022; Zhang et al., 2023a), current LLM based vision language models are much more powerful in terms of capabilities, there has not been much improvement in data utilization; they still rely on large-scale image-text pairs. Similarly, VLMs with dual encoders, as referred to in the first category above, also require large-scale image-text pair data. However, the data of the image-text pairs are gradually being exhausted (Villalobos et al., 2022). The total amount of a large number of cleaned image-text pairs is fixed, and the expansion of data scale can hardly keep up with the requirements of model scaling. Therefore, despite the remarkable advancements achieved in the VLM training paradigm, the utilization of large-scale image-text pairs data has encountered certain limitations. To overcome this issue, a potential solution lies in synthetic data.
Synthetic data. Synthetic data has been extensively utilized in large language models, yielding remarkable results (Zhang et al., 2024; Shao et al., 2024; Zhu et al., 2024; Yang et al., 2024). In addition, there is a growing presence of synthetic data within the multimodal domain. BLIP-2 (Li et al., 2023a) employs an image caption model to synthesize short captions to serve as substitutes for alt-text, and CapFusion (Yu et al., 2024) further enhances the quality of such data using large language models. However, the captions generated through this type of work tend to be rather simplistic, merely representing minor refinements of the original alt-text without adding substantial information value. Recap-DataComp-1B (Li et al., 2024b) aims to transform alt-text using vision language models, yet it encounters the issue that existing VLMs can produce severe hallucinations. Other works focused on dense captioning, such as Allava (Chen et al., 2024a) and Dense-Fusion-1M (Li et al., 2024c), are incapable of injecting rich knowledge in a targeted fashion and, on this basis, maintain relatively low levels of hallucination. In summary, knowledge injection and elimination of hallucinations are two key challenges in the realm of multimodal synthetic data. A significant amount of research has been dedicated to exploring the hallucination problem in VLMs, including aspects such as the causes underlying hallucinations and strategies for their elimination (Bai et al., 2024). However, a comprehensive method that can simultaneously ensure content diversity and maximally eliminate hallucinations in the synthetic visual caption task remains elusive.
Our approach has achieved substantial advances in both hallucination elimination and knowledge injection. Ultimately, its efficacy has been demonstrated in the training of VLMs, establishing our caption dataset as a viable alternative for the large-scale training of vision language models. This eliminates the constraints imposed by the data scale and avoids the challenges associated with data quality degradation resulting from the expansion of the data scale.
3 Methodology
3.1 Continuous DPO
Reinforcement learning has been reported to be an effective method of eliminating hallucinations in Large Language Models (LLMs) (Zhang et al., 2023b; Huang et al., 2025). Direct Preference Optimization (DPO) (Rafailov et al., 2023) emerges as a reinforcement learning based fine-tuning paradigm for LLM that directly optimizes preference data, effectively mitigating harmful content generation, such as hallucinations. Consequently, the application of DPO to mitigate hallucinations in image captions appears to be a natural and logical approach. The optimization objectives of the DPO algorithm are shown in Equation 1 and Equation 2. As noted in Rafailov et al. (2023), the loss function gradient $\mathcal{L}_{\text{DPO}}$ increases the probability of preferred completions $y_{w}$ while simultaneously reducing the probability of dispreferred completions $y_{l}$. Intuitively, by maximizing the probability of samples free of hallucinations and minimizing the probability of samples containing hallucinations, our recaption model can generate captions that are as free from uncertainty and potential hallucinations as possible.
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=$ (2)
$-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\underbrace{\sigma(\hat{r}_{\theta}(x,y_{l})-\hat{r}_{\theta}(x,y_{w}))}_{\text{higher weight when reward estimate is wrong}}\left[\underbrace{\nabla_{\theta}\log\pi(y_{w}\mid x)}_{\text{increase likelihood of }y_{w}}-\underbrace{\nabla_{\theta}\log\pi(y_{l}\mid x)}_{\text{decrease likelihood of }y_{l}}\right]\right],$
where $\hat{r}_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ is the reward implicitly defined by the language model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$.
However, in our task, the DPO algorithm does not demonstrate continuous performance improvement with respect to scale. Specifically, after initial rapid scaling up of performance, its performance plateaus, as illustrated in Figure 4. We turn our attention back to Equation2. Here, the examples are weighted according to how much higher the implicit reward model $\hat{r}_{\theta}(x,y)$ rates the dispreferred completions. As we continuously expand the scale of the DPO data, the implicit reward $\hat{r}_{\theta}(x,y)$, which serves as the optimization weight, does not change significantly. Consequently, after reaching a local optimum, the model ceases to optimize further. Therefore, we make slight modifications to the DPO algorithm and propose Continuous DPO (CDPO). Once the performance reaches a bottleneck phase, we switch to a different reference model, resample data using the current model, adjust the implicit reward model $\hat{r}_{\theta}(x,y)$ accordingly, and then re-optimize $\pi_{\theta}$. Furthermore, when CDPO training with preference data is performed, in order to prevent the sample length from serving as a learning signal (e.g., where longer captions may contain more hallucinations), we conduct an additional sequence-length balancing operation on the samples. This ensures that the average lengths of the positive and negative samples are close to each other.
3.2 Knowledge-enriching SFT
Previous recaption models (e.g.(Li et al., 2023a; Yu et al., 2024)) often produce generalized and imprecise descriptions, such as "a man performing on stage" in Figure 1. This type of descriptive information has limited informational value and offers little assistance in training vision language models. Hence, in our work, we place significant emphasis on the integration of valuable information. We believe that valuable information is mainly reflected in two aspects: one is the richness of descriptive details and the other is the knowledge content of the description. The former is easy to understand, while the latter is manifested in the supplementation of external knowledge related to the image. For example, in Figure 1, in the case of "oxalic acid", we hope to describe it as follows: The image shows the molecular chemical structure of oxalic acid, with the chemical formula $C_{2}H_{2}O_{4}$. The oxalic acid molecule consists of two carbon atoms, two hydrogen atoms, and four oxygen atoms. In the structural formula, the two carbon atoms are connected by single bonds, and each carbon atom is attached to a hydroxyl group ($OH$) and a carbonyl group ($C=O$), which are the typical structural features of oxalic acid.
Figure 2: A GPT-4o prompt designed to construct image captions rich in descriptive details. In the prompt, the red-highlighted portion lays stress on the methodology of knowledge injection. Furthermore, it guarantees that the ultimately generated caption incorporates a rich body of world knowledge via a two-step procedure.
By constructing SFT data, captions rich in descriptive details can be generated, enabling the model to learn to produce longer captions that contain as much specific information as possible through fine-tuning. Therefore, we carefully constructed a complex prompt and obtained training data using GPT-4o, see Section 3.3. Note that although GPT-4o can understand the requirements of the prompt, the output data still contain a substantial amount of hallucinations. As a result, we perform a manual review. Moreover, to incorporate knowledge into captions more effectively, we screened the original image-text pairs. In addition to general alt-text, we selected object-related data with a higher informational value [^1] In our study, we primarily regulate this aspect through the meticulous selection of alt-text data sources. For example, we deliberately choose a diverse range of encyclopedia data. This allows the descriptions in the captions to focus more on specific object descriptions, such as the "tragocephala pretiosa" in Figure 1. After reinforcement of the above two aspects, we construct knowledge-enhanced supervised fine-tuning data capable of supporting the training of image captions with rich content.
3.3 Training Pipeline
As mentioned above, our comprehensive training pipeline consists of two sequential stages: knowledge-enriching SFT followed by continuous DPO. In this section, we introduce these stages specifically. The general pipeline of our work is shown in Figure 3. We adopt Qwen2-VL-7B (Wang et al., 2024b) as the base model and use the LLaMA Factory (Zheng et al., 2024) platform to carry out the overall training process.
Figure 3: The illustration of our recaptioning pipeline
Knowledge-enriching SFT. We sampled raw image-text pairs from open-source multimodal datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), Datacamp (Gadre et al., 2023) Wukong (Gu et al., 2022), and Wikipeda (Srinivasan et al., 2021). Using the prompt shown in Figure 2, we employed GPT-4o to extract the image-alttext-caption data. Following a meticulous manual review process, $43,408$ data entries were retained. Subsequently, we performed LoRA fine-tuning on the Qwen2-VL-7B model. Specifically, the learning rate was set at $51.0e-$ 5, the global batch size was configured as $128$, and the number of training epochs was established at $10$. Through this process, we eventually acquired our Recaption-SFT model.
Continuous DPO. As outlined in Section 3.1, our methodology begins with training a DPO intermediate model. We first randomly sample $300,000$ image-text pairs and then employ the Recaption-SFT model for inference. To induce controlled stochasticity, we systematically modulate three sampling hyperparameters, TOPP=$1.0$, TOPK=$20$, and TEMPERATURE=$1.0$. Each image-text pair undergoes eight parallel inference iterations, thereby generating eight divergent candidates. We construct preference-ordered pairs by selecting the best and worst outputs from each set with a critic model  [^2] It is an in-house expert system that primarily makes point-by-point judgments by dissecting the details in captions and then assesses the hallucination level of the details. The main pipeline is similar to that of Section 4.1, but it does not use GPT-4o.
To eliminate length-induced training bias, we implement length-balanced sampling that retains $218k$ high-quality pairs, enforcing alignment of length distribution between positive and negative samples. Ultimately, the curated dataset was used for DPO training on the Recaption-SFT model, resulting in the DPO intermediate model.
Upon acquiring the intermediate model, we further sampled $200k$ data instances from the data sources and used this intermediate model to generate preference data with the same hyperparameters and the judge pipeline. Following the length-balancing strategy, we ultimately retained $139k$ data instances for continuous DPO training. All hyperparameters at different stages are listed in Table 2.
Stage-1-SFT Stage-2-DPO Stage-2.5-CDPO
Resolution # pixels [3136, 12845056] [3136, 12845056] [3136, 12845056]
Dataset # Samples 43K 218K 139K
Finetuning Type LoRA LoRA LoRA
Batch Size 128 64 64
Learning Rate $51e-$ 5 $65e-$ 6 $65e-$ 6
Learning Epoch 10 1 1
Table 2: Detailed configuration for each training stage of our recaption model. The table outlines the progression of vision parameters, dataset characteristics, model specifications and training hyperparameters.
4 Experiments
This section outlines our quality assessment criteria for recaption data and elaborates on our work from two aspects: the intrinsic quality of the dataset and its effectiveness in large-scale vision language model pre-training.
(a) Non-hallucination Rate
(b) Low-hallucination Rate
Figure 4: Performance of different DPO strategies. The horizontal axis represents the scale of the training data, and the vertical axis represents the proportion of hallucination-free captions on the validation set (in percentage). The CDPO strategy uses the same training data as DPO in the early stage of training and resamples $139k$ new data after training $218k$ data.
4.1 Evaluation Metrics
We believe that the quality measurement of recaption data encompasses two aspects. One is the assessment of harmful information, namely hallucinations. This information should be minimized as much as possible. The other aspect is the measurement of content richness, which can usually be verified by the effectiveness of model training. Therefore, in terms of the quality of the recaption data itself, we pay more attention to the indicators of hallucinations, while the measurement of its richness will be reflected through the training effect of large-scale vision language models in Section 4.3. We take advantage of the hallucination assessment method of CIEM (Hu et al., 2023) and leverage the existing large language model, GPT-4o, to decompose potential detailed elements within the caption. Subsequently, we use GPT-4o to make judgments about the presence of hallucinations. Finally, we count the number of hallucinations in a caption. We define the proportion of captions that do not contain hallucinatory details as the non-hallucination rate, and the proportion of captions that contain at most two hallucinatory details as the low-hallucination rate.
4.2 Performance of Recaption Model
We first report the evaluation results of our recaption model on open-source data, which come mainly from Wukong (Gu et al., 2022). As illustrated in Figure 4, the base model (Qwen2-VL-7B) exhibits a considerably severe level of hallucination. Specifically, its non-hallucination rate ($48.33\%$) and low-hallucination rate ($87.87\%$) are quite low. Following Supervised Fine-Tuning, the low-hallucination rate increases significantly. Subsequently, with the application of DPO, both the non-hallucination rate and the low-hallucination rate are substantially improved, by $34.03\%$ and $10.21\%$, respectively. Furthermore, we observe that DPO demonstrates a notable scaling law during the initial stage of the recaption task. However, the performance plateaus once the data size attains a certain scale. Our CDPO algorithm can further enhance both the non-hallucination rate (by $+3.4\%$) and the low-hallucination rate (by $+0.5\%$).
Meanwhile, we also use our model to infer on the same image-text pairs as other recaption works, e.g. Capfusion (Yu et al., 2024) and Recap-DataComp-1B (Li et al., 2024b), and report the results in Table 1. For the same original alt-text data, the recaption data generated by our model contain significantly more details than that of Capfusion and have much fewer hallucinations than Recap-DataComp-1B. Specifically, when dealing with the same data, compared to Capfusion, the number of details in the captions of our model increases by $267\%$, the average length of our captions is $3.5$ times longer than that of Capfusion, and the rate of hallucinatory details decreases by $8.6\%$. Moreover, compared to Recap-DataComp-1B, the non-hallucination rate increases by $43.6\%$. Note that since the length of captions in Capfusion is short, its non-hallucination rate is quite high. However, because of the lack of details, its value is limited.
Scientific QA OCR QA Benchmark QA Hallucination Tasks
Dataset MathVista MMMU AI2D ScienceQA OCRBench DocVQA TextVQA InfographicsVQA ChartQA MMBench MMStar RealworldQA MMVet HallusionBench None-Hallu. Rate Avg.
testmini val test test test val val val test (aug./hum.) en-dev test test test test
Alt-text(20M) 42.20 42.24 67.06 81.75 49.0 65.01 55.97 31.17 46.92 67.96 44.27 57.12 43.81 34.29 68.38 53.12
Recap-DataComp-1B(20M) 51.40 43.33 68.60 86.27 55.8 46.90 64.75 33.52 53.16 66.58 48.46 58.95 39.66 29.84 33.79 52.07
Hunyuan-Recap100M(20M) 52.50 44.27 73.59 85.82 57.3 66.27 72.59 42.47 57.00 72.77 50.13 60.26 46.02 35.45 73.87 59.35
Hunyuan-Recap100M 45.10 43.56 73.19 85.92 65.80 66.30 72.42 44.66 56.0 74.83 49.13 64.44 49.00 34.60 72.61 59.83
Alt-text(20M)$\ast$ 50.10 45.0 73.18 85.57 59.8 77.34 68.47 43.99 62.08 72.42 47.89 61.57 42.09 33.80 39.27 57.50
Recap-DataComp-1B(20M)$\ast$ 53.70 44.44 74.65 86.12 62.1 79.11 66.39 45.38 64.28 65.89 50.44 61.96 39.34 32.77 43.62 58.01
Hunyuan-Recap100M(20M)$\ast$ 56.20 44.00 77.91 89.74 65.9 83.06 71.37 50.32 68.96 72.34 54.66 65.49 43.17 36.38 71.86 63.42
[…]
Table 3: Experimental results on 15 vision-language tasks. Avg. is the average score of all vision-language tasks. The None-Hallu. Rate is a hallucination assessment metric introduced in Section 4.1. We randomly sampled 1000 pieces of DataCamp data to evaluate the level of hallucination. The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized. The $\ast$ indicates that a unified lite supervised fine-tuning (SFT) (Wang et al., 2023c; Li et al., 2024a) has been performed on the basis of pre-training. Note that our main purpose is to verify the impact of different caption datasets on the training of VLM models. Therefore, the above results are comparative experiments conducted under a fixed data setting (20M) without carefully supervised fine-tuning and reinforcement learning and are not comparable to SOTA models.
4.3 Performance of Recaption Data
We validate the effectiveness of Hunyuan-Recap100M in two scenarios, vision-language model pre-training and text-to-image generation. This validation is conducted primarily from the viewpoints of model training efficacy, hallucination levels, and knowledge coverage.
Vision-Language Model Pre-Training. We first validate the effectiveness of our data set in the pre-training of vision language models. The aim is to demonstrate that, compared to alt-text and other recaption data available on the market, Hunyuan-Recap100M can better assist in the training of vision language models and achieve better performance in both general open source multimodal tasks and in-house knowledge-related test sets. In the VLM framework, we select the LLaVA (Liu et al., 2023) architecture, with SigLIP (Zhai et al., 2023a) as visual encoder and Hunyuan-7B as the base large language model. To make a fair comparison of the impact of caption data, we sample $20M$ data points from alt-text/recaption data. Note that the alt-text group experiment uses the same images as Hunyuan-Recap100M, while Recap-Datacamp-1B uses randomly sampled data [^3] Since our data has undergone knowledge screening, its distribution is not exactly the same as that of DataCamp. As shown in Table 3, on 15 open source multimodal tasks, the VLM model trained with Hunyuan-Recap100M significantly outperforms those trained with alt-text or Recap-Datacamp-1B. Especially in hallucination-related tasks, Hunyuan-Recap100M shows a substantial lead over Recap-Datacamp-1B. Additionally, we evaluated the model’s convergence behavior through loss function monitoring. When using identical supervised fine-tuning datasets, models pre-trained with Hunyuan-Recap100M demonstrated markedly improved convergence properties.
In addition, to further measure the improvement in the perceptive level of the VLM caused by Hunyuan-Recap100M, we manually collected pictures with different visual characteristics from 20 domains, including animals, plants, clothing, landmarks, etc., totaling 1,197 pieces of data with manually annotated labels. We conducted an evaluation of the perceptive ability of the VLM in this set of data and the results are shown in Table 4. The VLM model trained with Hunyuan-Recap100M significantly outperforms other counterparts in 18 of the 20 categories. In general, it brings at least a $6.8\%$ improvement on average. After fully applying Hunyuan-Recap100M, the model has shown significant improvements in 17 of 20 perception tasks, with an additional average improvement of $9\%$.
Dataset Life Landmark Celebrity Food Cloth Good Game Traffic Art Festival Moive Icon Animal Plant Vegetables Sport Medical Military Tool History Avg.
Alt-text(20M) 34.1 35.9 23.4 22.4 26.2 22.7 24.5 34.5 23.3 27.1 25.5 27.5 54.2 27.3 22.2 65.5 32.6 29.4 20 27.3 30.0
Recap-DataComp-1B(20M) 13.6 19.7 18.9 4.5 29.9 23.2 21.6 33.9 12.2 10.0 12.7 17.0 49.2 13.6 13.1 57.9 19.6 19.6 30.4 17.6 21.8
Hunyuan-Recap100M(20M) 25.0 38.1 26.3 28.4 35.4 41.2 36.5 62.5 25.0 18.5 33.3 40.4 60.7 31.8 31.7 72.4 27.3 34.0 36.5 31.4 36.8
Hunyuan-Recap100M 40.0 44.3 39.0 34.4 53.8 38.1 51.9 63.6 27.5 35.1 31.5 51.0 66.1 50.7 50.8 79.3 46.3 39.6 36.4 38.2 45.8
Table 4: Experimental results on identification of 20 common vision object categories. The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized.
Text-to-Image Model Training We also verify the effectiveness of Hunyuan-Recap100M on the text-to-image generation task. We perform fine-tuning on the technical stack of Hunyuan-DiT (Li et al., 2024d) in order to demonstrate that our data can help generate specific visual images. As evidenced in Table 5, the publicly released Hunyuan-DiT model can significantly achieve a lower FID and a higher CLIP score on the MSCOCO dataset, as well as our internal perceptual test, when using our recaption data. Furthermore, after fine-tuning the data with Hunyuan-Recap100M, there is a further improvement in both the FID and the CLIP Score. This finding indicates that our data can offer considerable assistance in the realm of world perception and also enable text-to-image models to achieve better performance in terms of authenticity. We present realistic cases in Section B.
[…]
Dataset Row Caption Our Recaption Fine-tuning with Our Recaption Row Caption Our Recaption Fine-tuning with Our Recaption
perceptive Test Set 62.38 49.13 45.33 0.323 0.344 0.357
MSCOCO Test set 28.79 21.09 15.46 0.307 0.305 0.313
Table 5: Text-to-Image Evaluation of Hunyuan-DiT on our internal perceptual test and MSCOCO-30K, w/wo fine-tuned with our recaption data. Note that, due to computational resource limitations, we only sampled $2M$ data from Hunyuan-Recap100M for LoRA fine-tuning.
5 Conclusion and Limitation
5.1 Conclusion
In this work, we tackle the issue of synthesizing captions that are rich in knowledge and exhibit low levels of hallucination. We present a novel pipeline for high-quality synthetic captions, featuring two effective techniques: continuous DPO and knowledge-enriching SFT. We elucidate the beneficial role of DPO in alleviating hallucinations within image captions, along with its scaling law within a certain scope. Furthermore, we propose an approach to continually improve model performance once the scaling law ceases to be effective. Experimental results demonstrate that Hunyuan-Recap100M outperforms original alt-text and other synthetic captions in the pre-training of vision language models by a large margin. We will make Hunyuan-Recap100M publicly available to promote the development of multimodal related fields in the open source community.
Limitations.
Like most of the existing work on vision language models, the entire project consumed a lot with high electricity costs, although we only tested a $7B$ size model on $100M$ training data. Due to limited computing power, we did not scale up the use of recaption data to a large extent in this paper. We leave this work for future research. We also did not verify the effectiveness of the large-scale recaption data we generated for training text-to-image models from scratch. This will also be considered as future work.
Potential Risks.
The original image-text pairs are mainly derived from open source datasets. Although we have taken measures to detoxify to a great extent, there are still potential risks, especially in the field of image generation, such as fake portraits on social media (Hill and White, 2020), which is a common potential risk in image generation research.
Appendix A Loss of Vision-Language Model Training
To address the issue of incomparability of loss arising from diverse data distributions, we examine the convergence of pre-trained models that are trained on different datasets but are evaluated on the same supervised fine-tuning data, as illustrated in Figure 5. In particular, it can be observed that models trained with Hunyuan-Recap100M display a better convergence performance when dealing with the same data from downstream tasks. Specifically, these models exhibit lower loss values at both the starting point, with a loss value of $0.925$, and the convergence point, where the loss value is $0.48$.
Figure 5: Loss curve of vision-language model pre-trained with alt-text, Recap-DataComp-1B or Hunyuan-Recap100M. Note that all pre-trained models use the same amount of data (20M), and at the same time, the three curves are evaluated on exactly the same supervised fine-tuning data.
Appendix B Examples of Text-to-Image Generations
In this section, we present a series of examples of text-to-image generation in Figure 6. The synthetic captions corresponding to the text concepts in Figure 6 can be found in Table 6. From these examples, we can see that the HunyuanDiT model does not respond to a few concepts, such as abacus and Yueqin, and its depictions of some other concepts are also inaccurate, like jellyfish and acerola. After tuning with Hunyuan-Recap100M, the model’s understanding of these concepts becomes more precise, and its generated image style is also more realistic. The situations demonstrated in these cases are consistent with the reduction in FID shown in Table 5, highlighting the wide coverage of concepts perceived by our data. This is corroborated by the significant improvement in the performance of the VLM trained with Hunyuan-Recap100M on numerous object perception tasks, as presented in Table 4.
Text Image HunyuanDiT HunyuanDiT wRecap HunyuanDiT wFT
Jellyfish Refer to caption Refer to caption Refer to caption Refer to caption
Acerola Refer to caption Refer to caption Refer to caption Refer to caption
Abacus Refer to caption Refer to caption Refer to caption Refer to caption
Yueqin Refer to caption Refer to caption Refer to caption Refer to caption
Alpaca Refer to caption Refer to caption Refer to caption Refer to caption
The Grotto Refer to caption Refer to caption Refer to caption Refer to caption
Shanghai Tower Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Demonstration of text-to-image cases. Here, the original instruction for image generation is denoted as "Text", the original picture as "Image", the generation effect of the original open-source model as "HunyuanDiT", the generation effect of the original open-source model with our Synthetic texts as "HunyuanDiT wRecap", and the effect of the model fine-tuned with our recaption data as "HunyuanDiT wFT".
Text Our Recaption
Jellyfish The picture shows two jellyfishes. Their bodies are orange or light red with distinct stripes or lines. Their tentacles are long, slender, transparent or semi-transparent. The background is black, highlighting the colors of the jellyfish and the details of their tentacles. The jellyfish are swimming or floating in the water, and their tentacles are spread out in the water, forming a distinct linear effect.
Brazilian Acerola The image shows the branches of a tree laden with red and green fruits. The red fruits look ripe, while the green ones appear unripe. There are also green leaves on the branches. In the background, more green leaves and a few blurred flowers or small buds can be seen. The overall environment seems to be an outdoor natural setting.
Abacus There is a group of people sitting in front of a wooden table in the picture, using abacuses for learning or operation. There are several abacuses on the table. Wooden doors and windows can be seen in the background, indicating that this is an indoor environment, perhaps a classroom or a study place. People are wearing casual or semi-formal clothes, looking relaxed.
Yueqin This is a Yueqin, which has a circular resonator and four strings. There are multiple positions for pressing the strings on the body of the instrument, and there are four tuning pegs on the headstock. Overall, it exhibits the characteristics of a traditional musical instrument.
Alpaca There are two alpacas standing on the grass in the image, with blue sky and white clouds in the background. The alpaca on the left has lighter fur, while the one on the right has darker fur. The grass is green, and the postures of the alpacas suggest that they are either walking or standing.
The Grotto The Grotto is a church located in a natural environment. The facade of the church is a huge rock cave, covered with green vegetation above the cave, and surrounded by dense trees. Inside the church, there is an altar, above which a cross is hung, surrounded by lighting. In front of the church, there is a row of wooden benches for believers to sit and pray.
Shanghai Tower Shanghai Tower is a skyscraper with a spiral glass curtain wall design on its exterior. At the top, there is a circular observation deck, surrounded by other high-rise buildings. The overall architectural style is modern and futuristic.
Table 6: The original text of the image in Figure 6 and the synthetic caption after being processed by our recaption model.

### v2

@@
 Our continuous DPO methodology yields remarkable results in reducing hallucinations.
-Specifically, the non-hallucination caption rate on a held-out test set increases from $48.2\%$ to $77.9\%$ for a 7B-size model.
+Specifically, the non-hallucination caption rate on a held-out test set increases from $48.3\%$ to $77.9\%$ for a 7B-size model.
 2) Comprehensive empirical validation reveals that our synthetic captions confer superior pre-training advantages over their counterparts.
-Across 15 vision language tasks, the model trained with our data achieves a significant performance gain of at least $6.2\%$ compared to alt-text pairs and other previous work.
-In 20 common cognitive domains, the model trained with our data outperforms other counterparts by at least $6.8\%$.
+Across 15 vision language tasks, the model trained with our data achieves a significant performance gain of at least $6.2\%$ compared to identical images with alt-text.
+In 20 common cognitive domains, the model trained with our data outperforms the alt-text data by at least $7.5\%$.
 Meanwhile, it also offers considerable support in the text-to-image domain.
@@
 1 Introduction
-With the enormous power of large language models, remarkable performance gains have recently been achieved in a variety of tasks in natural language processing (NLP), computer vision (CV), and also in cross-modal fields (Brown et al., 2020; Chung et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a; Dosovitskiy et al., 2021; Alayrac et al., 2022; Liu et al., 2023; Achiam et al., 2023; Wang et al., 2024b; Hurst et al., 2024).
+With the enormous power of large language models, remarkable performance gains have recently been achieved in a variety of tasks in natural language processing (NLP), computer vision (CV), and also in cross-modal fields (Brown et al., 2020; Chung et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023; Dosovitskiy et al., 2021; Alayrac et al., 2022; Liu et al., 2023; Achiam et al., 2023; Wang et al., 2024; Hurst et al., 2024).
 The Large Language Model (LLM) and the Vision Language Model (VLM) are usually equipped with Transformer (Vaswani et al., 2017) as the backbone and then pre-trained with a tremendous amount of unlabeled data.
@@
 While text data has experienced exponential growth and fuels large language models, vision language model pre-training confronts a unique data bottleneck: high-quality text-image alignments remain scarce despite their critical role in training vision language models.
-Most traditional approaches to building large-scale multimodal datasets mainly rely on Web-based alt-text extraction (Sharma et al., 2018; Changpinyo et al., 2021; Ordonez et al., 2011; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Gadre et al., 2023; Liu et al., 2022; Gu et al., 2022).
-The web-based alt-text data typically undergoes a lengthy processing pipeline, involving stages such as text quality cleaning, image low-quality filtering, and cross-modal relevance validation, ultimately producing image-text pairs.
-Although significant data loss occurs at each processing stage and even with these efforts, the final output still suffers from inherent limitations including a low text informativeness density and partial relevance deficiencies between visual content and the corresponding descriptions.
-As the naturally existing image-text pair data is exhausted, and the quality of data produced by traditional image-text extraction methods drops significantly as the scale expands, the marginal effect they generate in the training of VLMs greatly diminishes.
-Therefore, new methods are needed to synthesize large-scale, high-quality multimodal data.
+Previous approaches to building large-scale multimodal datasets rely on Web-based alt-text extraction (Sharma et al., 2018; Changpinyo et al., 2021; Ordonez et al., 2011; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Gadre et al., 2023; Liu et al., 2022; Gu et al., 2022).
+The alt-text data undergo a lengthy processing pipeline, involving stages such as text cleaning, image filtering, and cross-modal relevance validation, ultimately producing image-text pairs.
+Although significant data loss occurs at each processing stage and even with these efforts, the final output still suffers from inherent limitations such as low informativeness and partial misalignment between visual content and the corresponding descriptions.
+As naturally occurring high-quality image-text pairs are increasingly exhausted, traditional extraction methods experience considerable declines in output quality when scaled up, leading to significantly diminishing marginal returns in the training of VLMs.
+Therefore, novel techniques are urgently needed to produce large-scale, high-quality multimodal datasets.
 Recently, several exploratory investigations have been carried out in multi-modal synthetic data, which have, to some extent, demonstrated their positive impact on various cross-modal tasks (Li et al., 2023a; Betker et al., 2023; Yu et al., 2024; Li et al., 2024b).
-However, there are significant defects in all of them, mainly reflected in the presence of many hallucinations or a relatively low level of information content.
+However, there are significant defects in all of them, mainly reflected in the presence of hallucinations and low information density.
 As shown in Figure 1, there are a large number of hallucinations (labeled in red) in the previous synthetic captions, and the captions are rather monotonous, lacking informative expressions.
@@
 How to establish an effective pipeline for generating low-hallucination synthetic captions?
-Large-scale synthetic captions inherently depend on robust tools, and vision language models emerge as a natural and fitting choice.
-However, hallucination is an inherent problem in the multimodal large language model (Bai et al., 2024).
-As shown in Figure 1, the data in the Recap-DataComp-1B (Li et al., 2024b) dataset contains a significant amount of hallucinatory information.
-We further demonstrate the harmfulness of a large amount of hallucinatory data in Section 4.
-The inevitable and most crucial issue in caption synthesis lies in resolving the hallucination problem.
-Existing work (Bai et al., 2024) has discussed the causes of hallucinations from the perspectives of data, model, training, and inference.
-During generation, as the length of the sequence grows, self-attention will focus more on previously generated text tokens, i.e., the attention to visual content is diluted(Wang et al., 2023b, a, 2024a; Huang et al., 2024).
-We agree that hallucinations in VLMs are mainly caused by uncertainty in decoding visual-related intentions.
-Therefore, we propose the use of Direct Preference Optimization (DPO) method to eliminate hallucinations in synthesized captions by increasing the confidences of real visual contents and suppressing the confidences of hallucinatory visual contents.
-Fortunately, this method has proven to be highly effective.
-Furthermore, in the synthetic caption task, we reveal that DPO adheres to a significant scaling law within a specific data size range.
-Upon saturation, resampling can further elevate the quality of synthetic captions.
-As presented in Table 1, when applied to the same dataset, our method significantly enhances the non-hallucination rate, elevating it from 29.7% to 73.3%, in contrast to the approach used in Recap-DataComp-1B (Li et al., 2024b).
-We elaborate on our hallucination evaluation metrics in Section 4.1 and the continuous DPO method in Section 3.
+VLMs enable efficient large-scale caption synthesis, but persistent hallucination effects constrain their reliability (Bai et al., 2024).
+As shown in Figure 1, Recap-DataComp-1B (Li et al., 2024b) contains a significant amount of hallucinatory information.
+Excess hallucinatory information in the training data is ultimately encoded in the final model, hindering the learning of correct patterns.
+The severity of this problem is quantitatively demonstrated in Section 4.
+Given the propensity for models trained on imperfect data to produce unfounded descriptions, effectively combating the generation of hallucinatory captions remains a paramount and often unavoidable challenge in caption synthesis.
+To address the crucial issue of hallucination in synthesized captions, we propose employing the Direct Preference Optimization (DPO) method.
+This approach aims to mitigate hallucinations by increasing the model’s confidence in generating descriptions of real visual content while simultaneously suppressing the confidence associated with hallucinatory visual elements.
+Furthermore, within the context of the synthetic caption task, we observe that DPO adheres to a significant scaling law within a specific data-size range.
+Notably, upon saturation of this scaling behavior, resampling of the data can further enhance the quality of the synthetic captions.
+As demonstrated in Table 1, applying our proposed method (more detailed in Section 3) to the same dataset yields a significant improvement in the non-hallucination rate (evaluated using metrics described in Section 4.1).
+Specifically, our approach increases the non-hallucination rate from 29.7% to 73.3%, representing a substantial enhancement compared to the baseline approach used in Recap-DataComp-1B (Li et al., 2024b).
 Dataset # Length # Details $\uparrow$ # None Hallucinations $\uparrow$ # Hallucination Rate $\downarrow$
@@
 Hunyuan-Recap100M 103.15 8.1 77.9% 4.2%
-Table 1: A comparison between existing synthetic caption datasets and our proposed dataset in terms of the number of hallucinations and the hallucination rate.
+Table 1: A comparison between existing synthetic caption datasets and our proposed dataset in terms of hallucinations.
 Here, "Length" represents the average length of the caption, "Details" represents the average number of visual details that appear in each caption, "None Hallucinations" indicates the proportion of non-hallucinatory data in the overall data, and "hallucination rate" represents the proportion of hallucinations among all details.
-A total of 1,000 data samples were taken, and our method reports the corresponding results on the respective data with our synthetic caption pipeline respectively.
+A total of 1,000 data samples were taken for validation in each dataset.
 How can more valuable information be injected into synthetic captions to increase the information density of synthetic data?
-The answer to this question can be rather straightforward: infuse as much knowledge as possible into the caption via a well-trained VLM.
-The injection of knowledge into synthetic captions can be derived from two primary sources.
-The first source is external text, such as alt-text, that is relevant to the image, while the second source stems from world knowledge encapsulated within a VLM.
-Usually, a VLM model tends to provide captions that are vague but correct.
-Therefore, we need to train a VLM model so that it can fully integrate the knowledge in text and its own world knowledge, ultimately generating synthetic captions with high informational value.
-Specifically, in our methodology, we employ a powerful VLM, namely GPT-4o, in conjunction with a meticulously crafted prompt that encompasses detailed requirements.
-This combination is utilized to generate labeled data, with a particular emphasis on extracting and focusing on the key information embedded within alt-text.
-For instance, we focus more on the name of a specific entity, using oxalic acid as an illustration here.
-After that, we perform manually post-processing on obvious erroneous data to ensure that the labeled data is accurate, contains the key information, and excludes the noise in the injected information.
-Finally, we train an open source VLM model, specifically Qwen2-VL-7B, using the training data.
-For details of the prompt and data processing, see Section 3.
-By comprehensively leveraging the aforementioned methods, we train a recaption model and utilize it to generate 100 million pieces of data.
-Our model has achieved a notable enhancement in the evaluation metrics for low-hallucination quantification.
-Moreover, we carried out a verification of the data’s validity.
-The results show that vision-language models pre-trained with Hunyuan-Recap100M exhibit substantially improved overall performance compared to those using alt-text data or conventional synthetic captions, particularly showing enhanced capabilities in perception tasks.
-Our contributions are as follows.
-(1) We tackle the challenge of constructing a synthetic caption pipeline capable of generating captions that are both low in hallucination and rich in knowledge, thereby alleviating the scarcity of high-quality multimodal data essential for vision-language model pre-training.
-(2) We propose a series of effective methods, knowledge-enriching SFT and continuous DPO, to eliminate hallucinations and inject knowledge into synthetic captions.
-Eventually, we trained a recaption model, whose proportion of hallucination-free results far exceeds that of the existing best models.
-The proportion of hallucination-free results has increased from $48.2\%$ to $77.9\%$.
-It also exhibits favorable performance in knowledge diversity, as reflected in a more accurate cognition of real-world entities.
-(3) We will release a high-quality synthetic caption dataset comprising $100M$ image-text pairs, Hunyuan-Recap100M, with the aim of fostering further advancements within the research community in areas like vision language model pre-training, text-to-image, and text-to-video generation.
+VLM outputs often prioritize visual saliency, resulting in captions that are generally correct but informatively sparse, lacking deeper context and specific details derived from external or implicit knowledge.
+To overcome this limitation and produce synthetic captions with high information density, our approach involves training a VLM to effectively synthesize both external textual knowledge (e.g.
+alt-text, metadata) and its inherent world knowledge.
+Specifically, we used a powerful VLM, GPT-4o, guided by a meticulously designed prompt, to create the initial set of labeled data.
+A key aspect of this data generation process is the focused extraction and integration of vital information embedded in external text sources, which serves as a readily available source of descriptive and often factual information about images.
+For example, particular attention is paid to identifying and incorporating specific entity names or critical attributes within the alt-text, illustrated here by the case of oxalic acid in Figure 1.
+Subsequently, the generated data undergoes a rigorous manual post-processing step to rectify obvious errors, confirm the inclusion of essential key information targeted during the generation phase, and effectively filter out noise or irrelevant details introduced during the initial knowledge injection.
+Finally, an open-source VLM, Qwen2-VL-7B, is fine-tuned on this carefully curated and information-rich dataset.
+Building upon the meticulously designed data generation and model training pipeline described previously, we developed and trained a specialized recaption model.
+We then proceeded to generate a large-scale synthetic dataset, Hunyuan-Recap100M, comprising 100 million image-text pairs.
+Compared to models pre-trained using raw alt-text data or conventional synthetic captions, VLMs trained with the Hunyuan-Recap100M dataset exhibit substantially improved overall performance.
+This significant performance gain, particularly pronounced in tasks requiring fine-grained visual perception, highlights the critical impact of training on synthetic data that not only offers high information density (as discussed previously), but also incorporates significantly reduced hallucinatory content, thereby enhancing the robustness and factual grounding of the downstream models.
+Empirical tests show models trained with our data gain at least 6.2% on 15 VL tasks and 6.8% on 20 cognitive domains compared to alternatives.
+This data also benefits text-to-image generation, reducing FID on validation benchmarks by up to 17.1.
+Our contributions are summarized as follows:
+1.
+High-Quality Synthetic Data Generation: We develop a novel pipeline leveraging specialized methods to generate synthetic captions that are both significantly low in hallucination (improving the hallucination-free rate from $48.3\%$ to a notable $77.9\%$) and rich in knowledge.
+2.
+Large-Scale Dataset and Demonstrated Impact: We construct and will release Hunyuan-Recap100M, a substantial 100 million image-text dataset generated by our pipeline.
+We demonstrate that pre-training vision-language models with Hunyuan-Recap100M leads to substantially improved overall performance compared to existing synthetic captions, particularly enhancing perceptual capabilities, thus providing a valuable resource to the community.
 2 Related Work
 Vision-Language Model.
-Vision Language Models (VLMs) can be classified into two distinct categories.
-The first category consists of models that rely on dual encoders, specifically a vision encoder and a text encoder.
-These models are typically trained in a contrastive fashion and exhibit remarkable proficiency in tasks such as image recognition.
-Prominent examples within this group include CLIP (Radford et al., 2021), FLIP (Li et al., 2023b) and SigLIP (Zhai et al., 2023b).
-The second category of methods focuses on aligning the visual modality with a Large Language Model (LLM).
-With the rapid development of LLM, the astonishing success of GPT (Brown et al., 2020; OpenAI, 2023), LLaMA (Touvron et al., 2023a, b; Grattafiori et al., 2024), and DeepSeek-LLM (Liu et al., 2024; Guo et al., 2025) has made the empowerment of large language models with visual perception an important and popular research direction.
-Flamingo (Alayrac et al., 2022) is the first large-scale vision language model, which aligns the visual encoder with a large language model.
-LLaVA (Liu et al., 2023) presents an instruction tuning methodology for training large multimodal models, thus charting the course for VLM training.
-Subsequent efforts have been made to refine the details of each module and continuously optimize the overall performance of the model (Wang et al., 2024b; Wu et al., 2024; McKinzie et al., 2024; Chen et al., 2024b; Liu et al., 2024).
-Although, compared with former multimodal models (Zeng et al., 2021; Singh et al., 2021; Diao et al., 2022; Wang et al., 2022; Zhang et al., 2023a), current LLM based vision language models are much more powerful in terms of capabilities, there has not been much improvement in data utilization; they still rely on large-scale image-text pairs.
-Similarly, VLMs with dual encoders, as referred to in the first category above, also require large-scale image-text pair data.
-However, the data of the image-text pairs are gradually being exhausted (Villalobos et al., 2022).
-The total amount of a large number of cleaned image-text pairs is fixed, and the expansion of data scale can hardly keep up with the requirements of model scaling.
-Therefore, despite the remarkable advancements achieved in the VLM training paradigm, the utilization of large-scale image-text pairs data has encountered certain limitations.
-To overcome this issue, a potential solution lies in synthetic data.
+Vision Language Models (VLMs) can be broadly classified into two categories.
+The first utilizes dual encoders, typically comprising a vision and a text encoder.
+These models are trained contrastively to learn aligned cross-modal representations and are highly effective in tasks such as image-text retrieval and zero-shot classification.
+Prominent examples include CLIP (Radford et al., 2021), FLIP (Li et al., 2023b), and SigLIP (Zhai et al., 2023b).
+The second category focuses on integrating visual information with a Large Language Model (LLM).
+Driven by the remarkable success of LLMs (Brown et al., 2020; OpenAI, 2023; Grattafiori et al., 2024; Guo et al., 2025), endowing these models with visual understanding has become a significant research direction.
+Flamingo (Alayrac et al., 2022) pioneered large-scale VLM research by aligning a visual encoder with an LLM.
+LLaVA (Liu et al., 2023) further established a methodology for training large multimodal models through instruction tuning.
+Subsequent efforts have continued to refine specific modules and optimize the overall performance of VLMs (Wang et al., 2024; Wu et al., 2024; McKinzie et al., 2024; Chen et al., 2024b; Liu et al., 2024).
+Although current LLM-based vision language models demonstrate significantly greater capabilities compared to earlier multimodal models (Zeng et al., 2021; Singh et al., 2021; Diao et al., 2022; Wang et al., 2022; Zhang et al., 2023a), they fundamentally share a core dependency with dual encoder VLMs (the first category discussed previously): the need for large-scale image-text pairs for training.
+This reliance presents a growing challenge, however, as the supply of high-quality image-text pair data is finite (Villalobos et al., 2022) and its expansion struggles to keep pace with the increasing scale of models.
+To overcome this data bottleneck, synthetic data offers a promising potential solution.
 Synthetic data.
-Synthetic data has been extensively utilized in large language models, yielding remarkable results (Zhang et al., 2024; Shao et al., 2024; Zhu et al., 2024; Yang et al., 2024).
-In addition, there is a growing presence of synthetic data within the multimodal domain.
-BLIP-2 (Li et al., 2023a) employs an image caption model to synthesize short captions to serve as substitutes for alt-text, and CapFusion (Yu et al., 2024) further enhances the quality of such data using large language models.
-However, the captions generated through this type of work tend to be rather simplistic, merely representing minor refinements of the original alt-text without adding substantial information value.
-Recap-DataComp-1B (Li et al., 2024b) aims to transform alt-text using vision language models, yet it encounters the issue that existing VLMs can produce severe hallucinations.
-Other works focused on dense captioning, such as Allava (Chen et al., 2024a) and Dense-Fusion-1M (Li et al., 2024c), are incapable of injecting rich knowledge in a targeted fashion and, on this basis, maintain relatively low levels of hallucination.
-In summary, knowledge injection and elimination of hallucinations are two key challenges in the realm of multimodal synthetic data.
-A significant amount of research has been dedicated to exploring the hallucination problem in VLMs, including aspects such as the causes underlying hallucinations and strategies for their elimination (Bai et al., 2024).
-However, a comprehensive method that can simultaneously ensure content diversity and maximally eliminate hallucinations in the synthetic visual caption task remains elusive.
-Our approach has achieved substantial advances in both hallucination elimination and knowledge injection.
-Ultimately, its efficacy has been demonstrated in the training of VLMs, establishing our caption dataset as a viable alternative for the large-scale training of vision language models.
-This eliminates the constraints imposed by the data scale and avoids the challenges associated with data quality degradation resulting from the expansion of the data scale.
+Building on its success in enhancing large language models (Zhang et al., 2024; Shao et al., 2024; Zhu et al., 2024; Yang et al., 2024), synthetic data generation is increasingly being explored within the multimodal domain.
+Some methods, such as BLIP-2 (Li et al., 2023a), utilize image captioning models to synthesize short captions as substitutes for alt-text, with CapFusion (Yu et al., 2024) leveraging LLMs for refinement.
+However, captions generated by these approaches often remain simplistic, providing minimal added information value beyond the original alt-text.
+Recap-DataComp-1B (Li et al., 2024b) attempts to transform alt-text using existing vision language models, but faces the significant challenge of hallucination inherent in current VLMs.
+Other works focusing on dense captioning, including Allava (Chen et al., 2024a) and Dense-Fusion-1M (Li et al., 2024c), generate localized descriptions, but the resulting datasets can also suffer from notable hallucinations.
+These limitations in existing synthetic VLM data highlight the need for more sophisticated generation techniques.
+In summary, generating high-quality multimodal synthetic data faces challenges related to knowledge enrichment and hallucination control.
+Although VLM hallucination is a studied topic (Bai et al., 2024), a comprehensive synthetic visual captioning method capable of simultaneously producing diverse and rich content with minimal hallucination is still needed.
+Our approach makes substantial progress in the generation of synthetic visual captions with improved knowledge injection and significantly reduced hallucination.
+This yields a caption dataset whose efficacy is demonstrated for large-scale VLM training, offering a viable alternative that overcomes the constraints of limited real data and the quality challenges of scaling traditional datasets.
 3 Methodology
+In this section, we present our synthetic data pipeline, including continuous DPO aimed at reducing hallucinations, knowledge-enriching SFT to enhance knowledge content, and details of the model training pipeline and hyper-parameters.
 3.1 Continuous DPO
-Reinforcement learning has been reported to be an effective method of eliminating hallucinations in Large Language Models (LLMs) (Zhang et al., 2023b; Huang et al., 2025).
+Reinforcement learning has been reported to be an effective method of mitigating hallucinations in large language models (Zhang et al., 2023b; Huang et al., 2025).
 Direct Preference Optimization (DPO) (Rafailov et al., 2023) emerges as a reinforcement learning based fine-tuning paradigm for LLM that directly optimizes preference data, effectively mitigating harmful content generation, such as hallucinations.
@@
 where $\hat{r}_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ is the reward implicitly defined by the language model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$.
-However, in our task, the DPO algorithm does not demonstrate continuous performance improvement with respect to scale.
-Specifically, after initial rapid scaling up of performance, its performance plateaus, as illustrated in Figure 4.
-We turn our attention back to Equation2.
-Here, the examples are weighted according to how much higher the implicit reward model $\hat{r}_{\theta}(x,y)$ rates the dispreferred completions.
-As we continuously expand the scale of the DPO data, the implicit reward $\hat{r}_{\theta}(x,y)$, which serves as the optimization weight, does not change significantly.
-Consequently, after reaching a local optimum, the model ceases to optimize further.
-Therefore, we make slight modifications to the DPO algorithm and propose Continuous DPO (CDPO).
-Once the performance reaches a bottleneck phase, we switch to a different reference model, resample data using the current model, adjust the implicit reward model $\hat{r}_{\theta}(x,y)$ accordingly, and then re-optimize $\pi_{\theta}$.
-Furthermore, when CDPO training with preference data is performed, in order to prevent the sample length from serving as a learning signal (e.g., where longer captions may contain more hallucinations), we conduct an additional sequence-length balancing operation on the samples.
-This ensures that the average lengths of the positive and negative samples are close to each other.
+However, in our task, standard DPO does not exhibit continuous performance improvement with respect to data scale.
+Specifically, after initial rapid performance gains, its performance plateaus, as illustrated in Figure 4.
+Examining the DPO objective (Equation 2), we see that the training examples are weighted based on the implicit reward derived from the reference model, which reflects the preference strength between chosen and rejected completions.
+As the scale of the DPO dataset increases, the implicit reward signal, derived from the fixed reference model, does not change significantly.
+Consequently, the optimization plateaus as the model converges to a local optimum defined by this relatively static signal.
+To address this limitation, we propose Continuous DPO (CDPO), a modification of the standard DPO algorithm.
+In CDPO, when performance plateaus, we iteratively update the reference model (e.g., using the current policy), resample preference data using the updated policy, which in turn yields a new implicit reward signal, and resume optimization.
+Furthermore, during CDPO training, to prevent the sample length from acting as an unintended learning signal (e.g.
+if longer completions are more likely to contain hallucinations), we apply a sequence-length balancing operation to the preference pairs.
+This ensures that the average lengths of the preferred and dispreferred completions within the dataset are closely matched.
 3.2 Knowledge-enriching SFT
-Previous recaption models (e.g.(Li et al., 2023a; Yu et al., 2024)) often produce generalized and imprecise descriptions, such as "a man performing on stage" in Figure 1.
-This type of descriptive information has limited informational value and offers little assistance in training vision language models.
-Hence, in our work, we place significant emphasis on the integration of valuable information.
-We believe that valuable information is mainly reflected in two aspects: one is the richness of descriptive details and the other is the knowledge content of the description.
-The former is easy to understand, while the latter is manifested in the supplementation of external knowledge related to the image.
-For example, in Figure 1, in the case of "oxalic acid", we hope to describe it as follows: The image shows the molecular chemical structure of oxalic acid, with the chemical formula $C_{2}H_{2}O_{4}$.
+Previous recaption models (e.g.
+(Li et al., 2023a; Yu et al., 2024)) often produce generalized descriptions lacking specificity, such as "a man performing on stage" in Figure 1.
+This type of descriptive information has limited informational value, thus offering little assistance in training vision language models.
+Hence, our work significantly emphasizes the integration of valuable information.
+We define valuable information as comprising two main aspects: the richness of descriptive details and the knowledge content of the description.
+The latter involves the incorporation of external knowledge related to the image.
+For example, considering "oxalic acid" in Figure 1, we aim for descriptions such as: The image shows the molecular chemical structure of oxalic acid, with the chemical formula $C_{2}H_{2}O_{4}$.
 The oxalic acid molecule consists of two carbon atoms, two hydrogen atoms, and four oxygen atoms.
 In the structural formula, the two carbon atoms are connected by single bonds, and each carbon atom is attached to a hydroxyl group ($OH$) and a carbonyl group ($C=O$), which are the typical structural features of oxalic acid.
-Figure 2: A GPT-4o prompt designed to construct image captions rich in descriptive details.
-In the prompt, the red-highlighted portion lays stress on the methodology of knowledge injection.
-Furthermore, it guarantees that the ultimately generated caption incorporates a rich body of world knowledge via a two-step procedure.
-By constructing SFT data, captions rich in descriptive details can be generated, enabling the model to learn to produce longer captions that contain as much specific information as possible through fine-tuning.
-Therefore, we carefully constructed a complex prompt and obtained training data using GPT-4o, see Section 3.3.
-Note that although GPT-4o can understand the requirements of the prompt, the output data still contain a substantial amount of hallucinations.
-As a result, we perform a manual review.
-Moreover, to incorporate knowledge into captions more effectively, we screened the original image-text pairs.
-In addition to general alt-text, we selected object-related data with a higher informational value [^1] In our study, we primarily regulate this aspect through the meticulous selection of alt-text data sources.
-For example, we deliberately choose a diverse range of encyclopedia data.
-This allows the descriptions in the captions to focus more on specific object descriptions, such as the "tragocephala pretiosa" in Figure 1.
-After reinforcement of the above two aspects, we construct knowledge-enhanced supervised fine-tuning data capable of supporting the training of image captions with rich content.
+Figure 2: A GPT-4o Prompt for Knowledge-Rich Captions.
+The portion highlighted in red within the prompt emphasizes the instructions for knowledge injection.
+This prompt utilizes a three-step procedure, which is intended to help ensure the inclusion of a rich body of world knowledge in the generated caption.
+By constructing SFT data with captions rich in descriptive details, we enable the model, through fine-tuning, to learn to produce detailed and specific descriptions.
+Therefore, we designed a detailed prompt and used GPT-4o to generate the initial training data.
+While GPT-4o understands the prompt requirements, the output data still contain a significant number of hallucinations.
+To address this, we conducted a manual review of the generated captions.
+By integrating these data processing steps, we constructed the knowledge-enhanced supervised fine-tuning data suitable for training models capable of generating rich image captions, such as the "arctolamia gestro" illustrated in Figure 1.
 3.3 Training Pipeline
@@
 The general pipeline of our work is shown in Figure 3.
-We adopt Qwen2-VL-7B (Wang et al., 2024b) as the base model and use the LLaMA Factory (Zheng et al., 2024) platform to carry out the overall training process.
-Figure 3: The illustration of our recaptioning pipeline
+We adopt Qwen2-VL-7B (Wang et al., 2024) as the base model and use the LLaMA Factory (Zheng et al., 2024) platform to carry out the overall training process.
+Figure 3: The illustration of our recaptioning pipeline.
+The pipeline comprises the following key stages: Initial Data Generation with GPT-4o, Manual Review, Knowledge-Enriching SFT, and Continuous DPO.
 Knowledge-enriching SFT.
-We sampled raw image-text pairs from open-source multimodal datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), Datacamp (Gadre et al., 2023) Wukong (Gu et al., 2022), and Wikipeda (Srinivasan et al., 2021).
-Using the prompt shown in Figure 2, we employed GPT-4o to extract the image-alttext-caption data.
+We sampled raw image-text pairs from open-source multimodal datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), DataComp (Gadre et al., 2023), Wukong (Gu et al., 2022), and Wikipedia (Srinivasan et al., 2021).
+Using the prompt in Figure 2, we employed GPT-4o to generate the initial recaption data.
 Following a meticulous manual review process, $43,408$ data entries were retained.
 Subsequently, we performed LoRA fine-tuning on the Qwen2-VL-7B model.
-Specifically, the learning rate was set at $51.0e-$ 5, the global batch size was configured as $128$, and the number of training epochs was established at $10$.
-Through this process, we eventually acquired our Recaption-SFT model.
+Specifically, training was conducted with a learning rate of $51.0e-$ 5, a global batch size of $128$, and for $10$ epochs.
+This process yielded our Recaption-SFT model.
 Continuous DPO.
-As outlined in Section 3.1, our methodology begins with training a DPO intermediate model.
-We first randomly sample $300,000$ image-text pairs and then employ the Recaption-SFT model for inference.
-To induce controlled stochasticity, we systematically modulate three sampling hyperparameters, TOPP=$1.0$, TOPK=$20$, and TEMPERATURE=$1.0$.
-Each image-text pair undergoes eight parallel inference iterations, thereby generating eight divergent candidates.
-We construct preference-ordered pairs by selecting the best and worst outputs from each set with a critic model  [^2] It is an in-house expert system that primarily makes point-by-point judgments by dissecting the details in captions and then assesses the hallucination level of the details.
-The main pipeline is similar to that of Section 4.1, but it does not use GPT-4o.
-To eliminate length-induced training bias, we implement length-balanced sampling that retains $218k$ high-quality pairs, enforcing alignment of length distribution between positive and negative samples.
-Ultimately, the curated dataset was used for DPO training on the Recaption-SFT model, resulting in the DPO intermediate model.
-Upon acquiring the intermediate model, we further sampled $200k$ data instances from the data sources and used this intermediate model to generate preference data with the same hyperparameters and the judge pipeline.
-Following the length-balancing strategy, we ultimately retained $139k$ data instances for continuous DPO training.
+As outlined in Section 3.1, our methodology begins with training the initial DPO model.
+We first randomly sample $300,000$ image-text pairs and then employ the SFT model for inference.
+To generate diverse candidates, we used the following sampling hyperparameters: TOPP=$1.0$, TOPK=$20$, and TEMPERATURE=$1.0$.
+Each image-text pair underwent eight parallel inference iterations, yielding eight divergent candidates.
+We then constructed preference pairs (chosen/rejected pairs) by selecting the best and worst outputs from each set using an internal critic model.
+[^1] This expert system evaluates caption quality by dissecting vision details in captions with an fine-tuned LLM (72B-sized) and assessing hallucination levels for each vision detail with a fine-tuned VLM (72B-sized)..
+To eliminate length-induced training bias, we implement length-balanced sampling that retains $218k$ high-quality pairs.
+This process ensures alignment in the length distribution between preferred and dispreferred completions.
+Subsequently, this curated data set was used for DPO training on the SFT model, yielding the initial DPO model (used as a reference model).
+Utilizing the trained initial DPO model, we sampled additional $200k$ image-text pairs from the data sources.
+These pairs were used to generate new preference data employing the model itself, the same sampling hyperparameters, and the critic model.
+Applying the length-balancing strategy to filter these generated pairs, we retained $139k$ preferred/dispreferred pairs for the subsequent CDPO training iteration.
 All hyperparameters at different stages are listed in Table 2.
@@
 4 Experiments
-This section outlines our quality assessment criteria for recaption data and elaborates on our work from two aspects: the intrinsic quality of the dataset and its effectiveness in large-scale vision language model pre-training.
+This section covers the quality assessment of our recaption data.
+We first outline the criteria used for this assessment, and then evaluate the data across two dimensions: its intrinsic quality and its impact on large-scale vision language model pre-training.
 (a) Non-hallucination Rate
@@
 Figure 4: Performance of different DPO strategies.
-The horizontal axis represents the scale of the training data, and the vertical axis represents the proportion of hallucination-free captions on the validation set (in percentage).
-The CDPO strategy uses the same training data as DPO in the early stage of training and resamples $139k$ new data after training $218k$ data.
+The horizontal axis represents the training data scale, while the vertical axis depicts the proportion of hallucination-free captions on the validation set (in percentage).
+Initially, the CDPO strategy utilizes the same training data as DPO.
+Upon reaching a training scale of $218k$ data points, it incorporates an additional $139k$ new data instances.
 4.1 Evaluation Metrics
-We believe that the quality measurement of recaption data encompasses two aspects.
-One is the assessment of harmful information, namely hallucinations.
-This information should be minimized as much as possible.
-The other aspect is the measurement of content richness, which can usually be verified by the effectiveness of model training.
-Therefore, in terms of the quality of the recaption data itself, we pay more attention to the indicators of hallucinations, while the measurement of its richness will be reflected through the training effect of large-scale vision language models in Section 4.3.
-We take advantage of the hallucination assessment method of CIEM (Hu et al., 2023) and leverage the existing large language model, GPT-4o, to decompose potential detailed elements within the caption.
-Subsequently, we use GPT-4o to make judgments about the presence of hallucinations.
-Finally, we count the number of hallucinations in a caption.
-We define the proportion of captions that do not contain hallucinatory details as the non-hallucination rate, and the proportion of captions that contain at most two hallucinatory details as the low-hallucination rate.
+The quality of recaption data is assessed along two key dimensions.
+One dimension concerns the presence of harmful information, particularly hallucinations, which must be minimized.
+The other is content richness, whose value is validated by its effectiveness in model training.
+Consequently, our primary quality control efforts focus on mitigating hallucinations.
+The richness of the data, conversely, is evaluated through the training performance of large-scale vision language models, as detailed in Section 4.3.
+Adopting the hallucination assessment method from CIEM (Hu et al., 2023), we utilize the large language model, GPT-4o, to parse each caption into its constituent visual details.
+GPT-4o then assesses each extracted detail for the presence of hallucinations.
+Following this detail-level judgment, we count the total number of hallucinatory details within a caption.
+Based on this count, we define two key metrics: the non-hallucination rate as the proportion of captions containing no hallucinatory details and the low-hallucination rate as the proportion of captions containing at most two hallucinatory details.
 4.2 Performance of Recaption Model
-We first report the evaluation results of our recaption model on open-source data, which come mainly from Wukong (Gu et al., 2022).
-As illustrated in Figure 4, the base model (Qwen2-VL-7B) exhibits a considerably severe level of hallucination.
-Specifically, its non-hallucination rate ($48.33\%$) and low-hallucination rate ($87.87\%$) are quite low.
-Following Supervised Fine-Tuning, the low-hallucination rate increases significantly.
-Subsequently, with the application of DPO, both the non-hallucination rate and the low-hallucination rate are substantially improved, by $34.03\%$ and $10.21\%$, respectively.
-Furthermore, we observe that DPO demonstrates a notable scaling law during the initial stage of the recaption task.
-However, the performance plateaus once the data size attains a certain scale.
-Our CDPO algorithm can further enhance both the non-hallucination rate (by $+3.4\%$) and the low-hallucination rate (by $+0.5\%$).
-Meanwhile, we also use our model to infer on the same image-text pairs as other recaption works, e.g.
-Capfusion (Yu et al., 2024) and Recap-DataComp-1B (Li et al., 2024b), and report the results in Table 1.
-For the same original alt-text data, the recaption data generated by our model contain significantly more details than that of Capfusion and have much fewer hallucinations than Recap-DataComp-1B.
-Specifically, when dealing with the same data, compared to Capfusion, the number of details in the captions of our model increases by $267\%$, the average length of our captions is $3.5$ times longer than that of Capfusion, and the rate of hallucinatory details decreases by $8.6\%$.
-Moreover, compared to Recap-DataComp-1B, the non-hallucination rate increases by $43.6\%$.
-Note that since the length of captions in Capfusion is short, its non-hallucination rate is quite high.
-However, because of the lack of details, its value is limited.
+To evaluate our model, we apply it to the identical image-text pairs used by other recaption methods, such as Capfusion (Yu et al., 2024) and Recap-DataComp-1B (Li et al., 2024b).
+The evaluation results are presented in Table 1.
+When generating data for these same input pairs, our model produces synthetic captions with significantly more details than those generated by Capfusion and substantially fewer hallucinations than those from Recap-DataComp-1B.
+Specifically, when evaluated on the same data, our model demonstrates substantial improvements over the baselines.
+Compared to Capfusion, our generated captions exhibit a $267\%$ increase in detail count, are $3.5$ times longer on average, and show a $8.6\%$ percentage point decrease in the rate of hallucinatory details.
+Relative to Recap-DataComp-1B, our model achieves a $43.6\%$ percentage point increase in the non-hallucination rate.
+It is important to note that while the brevity of Capfusion’s captions leads to a relatively high non-hallucination rate, the limited detail content restricts their overall value.
+We then report the evaluation results of our recaption model on another open source dataset, Wukong (Gu et al., 2022).
+As illustrated in Figure 4, the base model (Qwen2-VL-7B) exhibits a significant hallucination problem in this dataset.
+Specifically, its non-hallucination rate is $48.33\%$ and its low-hallucination rate is $87.87\%$, figures that indicate a substantial level of hallucinatory content.
+Following supervised fine-tuning, the low-hallucination rate is significantly increased.
+Subsequently, the application of DPO leads to substantial improvements in both the non-hallucination rate and the low-hallucination rate, increasing them $29.53\%$ and $10.21\%$, respectively.
+Furthermore, we observe that DPO exhibits a notable scaling law during the initial training stage for the recaption task; however, performance plateaus once the data size reaches a certain scale.
+Our CDPO algorithm provides further enhancement to both metrics, increasing the non-hallucination rate by $3.4\%$ and the low-hallucination rate by $0.5\%$.
 Scientific QA OCR QA Benchmark QA Hallucination Tasks
@@
 Recap-DataComp-1B(20M) 51.40 43.33 68.60 86.27 55.8 46.90 64.75 33.52 53.16 66.58 48.46 58.95 39.66 29.84 33.79 52.07
+Qwen2-VL-7B(20M) 54.40 42.33 73.77 87.31 61.3 62.88 71.9 40.92 56.68 69.24 48.86 59.74 48.20 31.84 65.13 58.31
 Hunyuan-Recap100M(20M) 52.50 44.27 73.59 85.82 57.3 66.27 72.59 42.47 57.00 72.77 50.13 60.26 46.02 35.45 73.87 59.35
@@
 Recap-DataComp-1B(20M)$\ast$ 53.70 44.44 74.65 86.12 62.1 79.11 66.39 45.38 64.28 65.89 50.44 61.96 39.34 32.77 43.62 58.01
+Qwen2-VL-7B(20M)$\ast$ 52 45.22 77.19 88.60 66.7 83.13 70.77 51.92 66.84 74.31 52.71 65.49 47.29 31.98 66.20 62.69
 Hunyuan-Recap100M(20M)$\ast$ 56.20 44.00 77.91 89.74 65.9 83.06 71.37 50.32 68.96 72.34 54.66 65.49 43.17 36.38 71.86 63.42
@@
 Avg.
-is the average score of all vision-language tasks.
-The None-Hallu.
-Rate is a hallucination assessment metric introduced in Section 4.1.
-We randomly sampled 1000 pieces of DataCamp data to evaluate the level of hallucination.
-The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized.
-The $\ast$ indicates that a unified lite supervised fine-tuning (SFT) (Wang et al., 2023c; Li et al., 2024a) has been performed on the basis of pre-training.
-Note that our main purpose is to verify the impact of different caption datasets on the training of VLM models.
-Therefore, the above results are comparative experiments conducted under a fixed data setting (20M) without carefully supervised fine-tuning and reinforcement learning and are not comparable to SOTA models.
+denotes the average score across all evaluated vision-language tasks.
+The Non-Hallucination Rate is a hallucination assessment metric defined in Section 4.1.
+For evaluation, we randomly sampled 1000 instances from the DataCamp dataset to quantify the non-hallucination rate.
+In the table, bold text indicates the optimal performance for a given metric under the $20M$ data setting, while underlined text signifies further enhancement achieved when utilizing the entire dataset.
+The $\ast$ symbol indicates that a light, unified supervised fine-tuning (SFT) (Wang et al., 2023; Li et al., 2024a) was applied after pre-training.
+It should be noted that our primary objective is to investigate the impact of different caption datasets on VLM model training.
+Consequently, these results represent comparative experiments conducted at a fixed data scale ($20M$).
+They were performed without employing advanced fine-tuning or reinforcement learning techniques often used to boost performance towards state-of-the-art, and thus are not directly comparable to SOTA models.
 4.3 Performance of Recaption Data
@@
 Vision-Language Model Pre-Training.
-We first validate the effectiveness of our data set in the pre-training of vision language models.
-The aim is to demonstrate that, compared to alt-text and other recaption data available on the market, Hunyuan-Recap100M can better assist in the training of vision language models and achieve better performance in both general open source multimodal tasks and in-house knowledge-related test sets.
-In the VLM framework, we select the LLaVA (Liu et al., 2023) architecture, with SigLIP (Zhai et al., 2023a) as visual encoder and Hunyuan-7B as the base large language model.
-To make a fair comparison of the impact of caption data, we sample $20M$ data points from alt-text/recaption data.
-Note that the alt-text group experiment uses the same images as Hunyuan-Recap100M, while Recap-Datacamp-1B uses randomly sampled data [^3] Since our data has undergone knowledge screening, its distribution is not exactly the same as that of DataCamp.
-As shown in Table 3, on 15 open source multimodal tasks, the VLM model trained with Hunyuan-Recap100M significantly outperforms those trained with alt-text or Recap-Datacamp-1B.
-Especially in hallucination-related tasks, Hunyuan-Recap100M shows a substantial lead over Recap-Datacamp-1B.
-Additionally, we evaluated the model’s convergence behavior through loss function monitoring.
-When using identical supervised fine-tuning datasets, models pre-trained with Hunyuan-Recap100M demonstrated markedly improved convergence properties.
-In addition, to further measure the improvement in the perceptive level of the VLM caused by Hunyuan-Recap100M, we manually collected pictures with different visual characteristics from 20 domains, including animals, plants, clothing, landmarks, etc., totaling 1,197 pieces of data with manually annotated labels.
-We conducted an evaluation of the perceptive ability of the VLM in this set of data and the results are shown in Table 4.
-The VLM model trained with Hunyuan-Recap100M significantly outperforms other counterparts in 18 of the 20 categories.
-In general, it brings at least a $6.8\%$ improvement on average.
-After fully applying Hunyuan-Recap100M, the model has shown significant improvements in 17 of 20 perception tasks, with an additional average improvement of $9\%$.
+We first validate the effectiveness of our dataset for the pre-training of vision language models.
+This aims to demonstrate that, compared to existing alt-text and other recaption datasets, Hunyuan-Recap100M can facilitate more effective training of vision language models, leading to superior performance on both general open-source multimodal tasks and our internal knowledge-related test sets.
+For the VLM framework, we adopt the LLaVA (Liu et al., 2023) architecture, comprising SigLIP (Zhai et al., 2023a) as visual encoder and Hunyuan-7B (Hunyuan, 2025) as the base large language model.
+To ensure a fair comparison of the influence of caption datasets, we sample $20M$ data points from each of the alt-text and recaption datasets.
+Note that the experiments that use the alt-text data utilize the same image set as our Hunyuan-Recap100M dataset.
+In contrast, the Recap-Datacamp-1B data uses a randomly sampled subset, implying a different image set distribution compared to the alt-text and Hunyuan-Recap100M data.
+[^2] Hunyuan-Recap100M dataset is compiled from multiple open-source datasets, resulting in a distribution that is not exactly the same as that of DataCamp.
+In Table 3, "Qwen2-VL-7B(20M)" refers to the model trained on the $20M$ recaption data generated by the unoptimized Qwen2-VL-7B model, which contains more hallucinations.
+On 15 open-source multimodal tasks, the VLM model pre-trained with Hunyuan-Recap100M data significantly outperforms models pre-trained with alt-text or other synthetic captions.
+This superiority is particularly evident in hallucination-related tasks, where training with Hunyuan-Recap100M data achieves a substantial lead over training with other counterparts.
+Additionally, we monitored the loss function to evaluate the model convergence behavior in the appendix A.
+When subjected to identical supervised fine-tuning datasets, models pre-trained with Hunyuan-Recap100M demonstrated markedly improved convergence properties.
+Furthermore, to quantify the enhancement in the visual perception capabilities of VLM attributable to Hunyuan-Recap100M, we manually collected $1,197$ images with diverse visual characteristics across 20 domains, including animals, plants, clothing, and landmarks, among others.
+These images are accompanied by manually annotated labels for evaluation.
+We assessed the VLM’s perceptive ability on this dataset, and the results are presented in Table 4.
+When trained with a $20M$ subset of Hunyuan-Recap100M data, the VLM model significantly outperforms other counterparts in 14 of the 20 categories, demonstrating an average improvement of at least $3.8\%$ compared to models trained on other $20M$ datasets.
+Upon utilizing the full Hunyuan-Recap100M dataset for training, the model shows significant improvements in 18 of the 20 perception tasks, with an overall average improvement reaching $9\%$.
 Dataset Life Landmark Celebrity Food Cloth Good Game Traffic Art Festival Moive Icon Animal Plant Vegetables Sport Medical Military Tool History Avg.
-Alt-text(20M) 34.1 35.9 23.4 22.4 26.2 22.7 24.5 34.5 23.3 27.1 25.5 27.5 54.2 27.3 22.2 65.5 32.6 29.4 20 27.3 30.0
-Recap-DataComp-1B(20M) 13.6 19.7 18.9 4.5 29.9 23.2 21.6 33.9 12.2 10.0 12.7 17.0 49.2 13.6 13.1 57.9 19.6 19.6 30.4 17.6 21.8
-Hunyuan-Recap100M(20M) 25.0 38.1 26.3 28.4 35.4 41.2 36.5 62.5 25.0 18.5 33.3 40.4 60.7 31.8 31.7 72.4 27.3 34.0 36.5 31.4 36.8
-Hunyuan-Recap100M 40.0 44.3 39.0 34.4 53.8 38.1 51.9 63.6 27.5 35.1 31.5 51.0 66.1 50.7 50.8 79.3 46.3 39.6 36.4 38.2 45.8
-Table 4: Experimental results on identification of 20 common vision object categories.
-The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized.
-Text-to-Image Model Training We also verify the effectiveness of Hunyuan-Recap100M on the text-to-image generation task.
-We perform fine-tuning on the technical stack of Hunyuan-DiT (Li et al., 2024d) in order to demonstrate that our data can help generate specific visual images.
-As evidenced in Table 5, the publicly released Hunyuan-DiT model can significantly achieve a lower FID and a higher CLIP score on the MSCOCO dataset, as well as our internal perceptual test, when using our recaption data.
-Furthermore, after fine-tuning the data with Hunyuan-Recap100M, there is a further improvement in both the FID and the CLIP Score.
-This finding indicates that our data can offer considerable assistance in the realm of world perception and also enable text-to-image models to achieve better performance in terms of authenticity.
-We present realistic cases in Section B.
+Alt-text(20M) 54.7 36.83 22.67 24.13 29.93 37.07 25.30 50.57 29.47 26.73 17.10 41.80 60.57 37.77 24.73 67.77 41.30 50.30 28.17 31.00 36.33
+Recap-DataComp-1B(20M) 32.57 26.60 20.87 21.53 34.47 30.07 24.83 46.57 19.47 10.20 14.73 25.50 54.57 28.57 26.93 60.67 24.63 36.13 37.60 21.93 29.67
+Qwen2-VL-7B(20M) 45.83 35.60 34.47 31.90 49.77 45.73 34.80 66.10 36.30 23.83 21.40 40.60 54.30 32.17 29.43 69.83 39.73 43.87 28.90 36.53 39.97
+Hunyuan-Recap100M(20M) 42.80 45.30 27.10 32.03 43.90 54.43 34.80 68.23 37.83 28.73 28.13 52.70 61.83 44.10 38.27 74.30 34.00 45.60 45.13 35.23 43.80
+Hunyuan-Recap100M 53.43 52.10 39.03 46.07 58.70 55.40 52.53 74.57 44.57 40.47 33.13 67.10 75.27 62.60 54.07 83.87 41.50 50.87 41.63 43.53 53.67
+Table 4: Evaluation results on the identification of 20 common visual object categories.
+Text in bold indicates the optimal performance for a given category under the $20M$ data setting, while underlined text signifies further performance enhancement achieved when utilizing the entire dataset.
+All reported scores were generated by GPT-4o on May 6, 2025.
+To mitigate scoring fluctuations, we conducted three independent scoring iterations and used the mean value as the final score.
+Text-to-Image Model Training We also evaluate the effectiveness of Hunyuan-Recap100M on the text-to-image generation task.
+By fine-tuning the Hunyuan-DiT (Li et al., 2024d), we aim to demonstrate that our data can improve visual generation capabilities.
+As evidenced in Table 5, models fine-tuned from the publicly released Hunyuan-DiT using our recaption data significantly achieve a lower FID and a higher CLIP score on the MSCOCO dataset and an internal perceptual test set.
+These findings indicate that our data offer considerable assistance in enhancing the model’s understanding of visual concepts and enabling text-to-image models to achieve better performance in terms of authenticity.
+Realistic cases are presented in Section B.
 […]
@@
 MSCOCO Test set 28.79 21.09 15.46 0.307 0.305 0.313
-Table 5: Text-to-Image Evaluation of Hunyuan-DiT on our internal perceptual test and MSCOCO-30K, w/wo fine-tuned with our recaption data.
-Note that, due to computational resource limitations, we only sampled $2M$ data from Hunyuan-Recap100M for LoRA fine-tuning.
+Table 5: Text-to-Image Evaluation of the Hunyuan-DiT model on an internal perceptual test set and MSCOCO-30K, comparing performance with and without fine-tuning using our recaption data.
+It should be noted that, due to computational resource limitations, we only sampled $2M$ data points from the Hunyuan-Recap100M dataset for LoRA fine-tuning.
 5 Conclusion and Limitation
-5.1 Conclusion
-In this work, we tackle the issue of synthesizing captions that are rich in knowledge and exhibit low levels of hallucination.
-We present a novel pipeline for high-quality synthetic captions, featuring two effective techniques: continuous DPO and knowledge-enriching SFT.
-We elucidate the beneficial role of DPO in alleviating hallucinations within image captions, along with its scaling law within a certain scope.
-Furthermore, we propose an approach to continually improve model performance once the scaling law ceases to be effective.
-Experimental results demonstrate that Hunyuan-Recap100M outperforms original alt-text and other synthetic captions in the pre-training of vision language models by a large margin.
-We will make Hunyuan-Recap100M publicly available to promote the development of multimodal related fields in the open source community.
-Limitations.
-Like most of the existing work on vision language models, the entire project consumed a lot with high electricity costs, although we only tested a $7B$ size model on $100M$ training data.
-Due to limited computing power, we did not scale up the use of recaption data to a large extent in this paper.
-We leave this work for future research.
-We also did not verify the effectiveness of the large-scale recaption data we generated for training text-to-image models from scratch.
-This will also be considered as future work.
-Potential Risks.
-The original image-text pairs are mainly derived from open source datasets.
-Although we have taken measures to detoxify to a great extent, there are still potential risks, especially in the field of image generation, such as fake portraits on social media (Hill and White, 2020), which is a common potential risk in image generation research.
+This work addresses the challenge of synthesizing low-hallucination, knowledge-rich image captions.
+We introduce a novel pipeline for generating such high-quality synthetic captions employing two key techniques: continuous DPO (CDPO) and knowledge-enriching SFT.
+We demonstrate that DPO effectively mitigates hallucinations in captions and analyze its scaling behavior at certain data scales.
+Furthermore, we present the CDPO approach designed to sustain performance improvements beyond the point where DPO’s scaling plateaus.
+Experimental results show that Hunyuan-Recap100M data significantly outperform original alt-text and other synthetic caption datasets for pre-training vision language models.
+To foster the development of multimodal research within the open source community, we will make Hunyuan-Recap100M publicly available.
+See Section 6 and Section 7 for details on limitations and risks.
+6 Limitations
+Training large vision-language models is inherently computationally intensive and resource-demanding.
+Specifically, pre-training the 7B-sized VLM model using the complete Hunyuan-Recap100M dataset required substantial $15,360$ H20 GPU compute hours.
+Due to limited computational resources, this work does not explore the full potential of scaling up the use of recaption data.
+Furthermore, we did not evaluate the effectiveness of using our generated large-scale recaption data to train text-to-image models from scratch.
+Addressing these limitations by conducting experiments with larger data scales and exploring T2I training from scratch are important avenues for future research.
+7 Potential Risks
+The original image-text pairs are primarily derived from open-source datasets.
+While we have implemented substantial measures to filter out undesirable content, potential risks remain.
+These risks are particularly salient in the field of image generation, exemplified by issues such as the creation of fake portraits for social media (Hill and White, 2020), which is a recognized challenge in this research area.
 Appendix A Loss of Vision-Language Model Training
-To address the issue of incomparability of loss arising from diverse data distributions, we examine the convergence of pre-trained models that are trained on different datasets but are evaluated on the same supervised fine-tuning data, as illustrated in Figure 5.
-In particular, it can be observed that models trained with Hunyuan-Recap100M display a better convergence performance when dealing with the same data from downstream tasks.
-Specifically, these models exhibit lower loss values at both the starting point, with a loss value of $0.925$, and the convergence point, where the loss value is $0.48$.
-Figure 5: Loss curve of vision-language model pre-trained with alt-text, Recap-DataComp-1B or Hunyuan-Recap100M.
-Note that all pre-trained models use the same amount of data (20M), and at the same time, the three curves are evaluated on exactly the same supervised fine-tuning data.
+To address the issue of loss value incomparability arising from diverse pre-training data distributions, we examine the convergence behavior of pre-trained models when fine-tuned on a common supervised dataset [Wang et al., 2023, Li et al., 2024a], as illustrated in Figure 5.
+Figure 5 shows that models trained with Hunyuan-Recap100M data exhibit superior convergence.
+Quantitatively, these models converge to a lower final loss of $0.4796$ compared to models pre-trained on other datasets.
+Figure 5: Loss curve of vision-language model pre-trained with alt-text, Recap-DataComp-1B, Qwen2-VL-Zeroshot or Hunyuan-Recap100M.
+To ensure a fair comparison, all pre-trained models analyzed here utilized the same amount of data ($20M$) for pre-training.
+Furthermore, the four convergence curves presented are obtained by fine-tuning these models on an identical supervised dataset.
 Appendix B Examples of Text-to-Image Generations
-In this section, we present a series of examples of text-to-image generation in Figure 6.
-The synthetic captions corresponding to the text concepts in Figure 6 can be found in Table 6.
-From these examples, we can see that the HunyuanDiT model does not respond to a few concepts, such as abacus and Yueqin, and its depictions of some other concepts are also inaccurate, like jellyfish and acerola.
-After tuning with Hunyuan-Recap100M, the model’s understanding of these concepts becomes more precise, and its generated image style is also more realistic.
-The situations demonstrated in these cases are consistent with the reduction in FID shown in Table 5, highlighting the wide coverage of concepts perceived by our data.
-This is corroborated by the significant improvement in the performance of the VLM trained with Hunyuan-Recap100M on numerous object perception tasks, as presented in Table 4.
-Text Image HunyuanDiT HunyuanDiT wRecap HunyuanDiT wFT
+This section presents examples of text-to-image generation in Figure 6; the corresponding text prompts are provided in Table 6.
+Observation of these examples reveals that the original Hunyuan-DiT model struggles to accurately render certain concepts (e.g.
+abacus, Yueqin) and produces less accurate depictions for others (e.g.
+jellyfish, acerola).
+However, after fine-tuning with the Hunyuan-Recap100M data, the model’s understanding of these concepts becomes more precise, resulting in generated images with a more realistic style.
+These qualitative improvements are consistent with the quantitative results shown in Table 5, such as the observed reduction in FID.
+This consistency highlights how the comprehensive concept coverage in our data enhances the model’s perceptual fidelity, which is further corroborated by the significant performance improvement of the VLM trained with Hunyuan-Recap100M on numerous object perception tasks (Table 4).
+Text Image Hunyuan-DiT Hunyuan-DiT wRecap Hunyuan-DiT wFT
 Jellyfish Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Figure 6: Demonstration of text-to-image cases.
-Here, the original instruction for image generation is denoted as "Text", the original picture as "Image", the generation effect of the original open-source model as "HunyuanDiT", the generation effect of the original open-source model with our Synthetic texts as "HunyuanDiT wRecap", and the effect of the model fine-tuned with our recaption data as "HunyuanDiT wFT".
+Here, the original instruction for image generation is denoted as "Text", the original picture as "Image", the generation effect of the original open-source model as "Hunyuan-DiT", the generation effect of the original open-source model with our Synthetic texts as "Hunyuan-DiT wRecap", and the effect of the model fine-tuned with our recaption data as "Hunyuan-DiT wFT".
 Text Our Recaption